# Accelerating Data Chunking in Deduplication Systems using Vector Instructions

### v1

@@ Abstract.
Content-defined Chunking (CDC) algorithms dictate the overall space savings that deduplication systems achieve. However, due to their need to scan each file in its entirety, they are slow and often the main performance bottleneck within data deduplication. We present VectorCDC, a method to accelerate hashless CDC algorithms using vector CPU instructions, such as SSE / AVX. Our evaluation shows that VectorCDC is effective on Intel, AMD, ARM, and IBM CPUs, achieving $8.35\times-26.2\times$ higher throughput than existing vector-accelerated techniques without affecting the deduplication space savings.
Content-defined chunking, SSE/AVX instructions, AVX-512, ARM NEON, IBM VSX
@@ 1. Introduction
The amount of data generated and stored on the Internet is growing at an exponential rate (statista2024, ), and is expected to exceed 180 zettabytes per year in 2025. Storage capacity alone is not well positioned to handle this data influx, with the total installed storage capacity in 2020 only being 6.7 zettabytes (statista2024, ). Cloud storage providers instead support this data growth using alternatives such as novel storage paradigms (software_defined_storage, ; raid, ), distributed file systems (hadoop, ; ceph, ) and caches (memcached, ; facebook_tao, ), mechanisms such as data deduplication (dedup_intro, ; comprehensive_dedup, ), alongside additional investments in data protection (data_security, ).
@@
Previous studies by Microsoft (primary_dedup, ) and EMC (backup_workload_charz, ) show that a large amount of redundancy exists in the data stored on the cloud, especially in file system backups (primary_dedup, ; dedup_intro, ), virtual machine backups (backup_workload_charz, ; wandeltacomp, ) and shared documents (documentduplication_2011, ). The redundancy ratios in these workloads are significant, ranging from $50\%$ to $75\%$. Mechanisms such as data deduplication (dedup_intro, ) and compression (wandeltacomp, ) are used to conserve storage space by identifying redundant data, eliminating it, and minimizing its storage impact, thereby improving efficiency and reducing costs.
@@
Data deduplication consists of four phases (comprehensive_dedup, ); Data Chunking, Chunk Fingerprinting, Metadata Creation, and Data Storage. Data chunking and chunk fingerprinting are the most computationally intensive (dedup_intro, ; dedupbench, ) of these. While chunk fingerprinting has received significant optimization attention, with faster hashing algorithms (tunable_encrypted_dedup, ; scalable_incremental_checkpt, ) and GPUs (gpu_dedup, ; storegpu, ), data chunking optimizations have not kept up (§3.1).
@@
In the data chunking phase, the incoming data is divided into small chunks, typically of size 1641-64 KB. Numerous data chunking algorithms exist in current literature (ae, ; fastcdc, ; gear_hash, ; lbfs, ; ram, ; tttd, ) and can be broadly classified into hash-based and hashless algorithms (dedupbench, ). As chunking occurs whenever new data is uploaded, this phase is on the critical path, and directly impacts system performance.
@@
Previous efforts explored accelerating chunking through using vector instructions. SS-CDC (sscdc, ) uses vector instructions to accelerate hash-based data chunking algorithms, such as Rabin-Karp chunking (lbfs, ) and Gear-based chunking (gear_hash, ). Unfortunately, this approach only leads to modest improvements in chunking throughput, up to $3.13\times$, as shown in §3.2. Parallelizing hash-based chunking using vector instructions is complicated because these algorithms use the rolling hash of a sliding window of bytes to detect boundaries, inherently creating a computational dependency between adjacent bytes. Consequently, SS-CDC processes different regions of the data in parallel. This approach relies on slow scatter / gather vector instructions, limiting its performance (§3.2).
We posit that hashless chunking algorithms are better candidates for vector acceleration. Although they achieve slightly lower space savings compared to their hash-based counterparts (§6), they are up $2\times$ faster and use simple mathematical operations (e.g., finding a maximum value) that can be accelerated more efficiently using vector instructions.
@@
We analyzed state-of-the-art hashless algorithms to understand their design and identify opportunities to leverage vector instructions. We identified that all state-of-the-art hashless algorithms consist of two processing patterns. The first pattern involves finding local minima or maxima in a data region, which we call Extreme Byte Search, and the second pattern involves scanning a range of bytes to find values that are greater or lesser than a target value, which we call Range Scan. We found that, unlike rolling hash functions, these patterns can be efficiently accelerated using vector instructions.
@@
We implemented VectorCDC using five different vector instruction sets: SSE-128, AVX-256, and AVX-512 on Intel / AMD CPUs; NEON-128 on ARM CPUs; and VSX-128 on IBM Power CPUs. We used VectorCDC to accelerate three state-of-the-art hashless chunking algorithms. We implemented VRAM, VAE, and VMAXP, which are accelerated versions of the RAM (ram, ), AE (ae, ), and MAXP (maxp, ) algorithms, respectively. We compared the performance of our accelerated algorithms against state-of-the-art hash-based algorithms, hashless algorithms, and SS-CDC (sscdc, ) accelerated algorithms using 10 diverse datasets.
Our evaluation (§6) shows that VectorCDC achieves $8.35\times$$26.2\times$ higher chunking throughput than SS-CDC’s approach, without affecting the space savings achieved by hashless algorithms. VectorCDC-based algorithms achieve up to $5.51\times$$17.6\times$ higher throughput compared to their unaccelerated counterparts. Furthermore, algorithms accelerated using VectorCDC retain their performance advantage with all five vector instruction sets. We have made our code publicly available by integrating it with DedupBench [^1] https://github.com/UWASL/dedup-bench (dedupbench, ). Due to the large sizes of our datasets (§6), we were unable to release them all. We have instead provided detailed descriptions to facilitate their recreation for future experiments, similar to previous literature (fastcdc, ; ae, ; ram, ). In addition, to facilitate reproducing our results, we have publicly released one of our datasets (DEB) on Kaggle [^2] https://www.kaggle.com/datasets/sreeharshau/vm-deb-fast25 (deb_dataset, ).
The rest of this paper is organized as follows: we discuss relevant background about deduplication and vector instructions in Section 2. Section 3 motivates our work by discussing deduplication performance bottlenecks and the inefficiencies encountered by previous work when accelerating hash-based CDC algorithms. Section 4 outlines VectorCDC’s design while Section 6 details our evaluation efforts on diverse datasets. We discuss related work in Section 7 and conclude our paper in Section 8.
@@ 2. Background
Data deduplication consists of four phases (comprehensive_dedup, ):
@@
  • Fingerprinting and Comparison: Chunks are hashed using a collision-resistant hashing algorithm such as MurmurHash3 (tunable_encrypted_dedup, ) or SHA-256 (sha256, ) to generate fingerprints. Fingerprints are compared against those previously seen to identify duplicate chunks.
@@
Data chunking and fingerprinting are typically the most computationally intensive phases in deduplication (dedup_intro, ). While fingerprinting has been accelerated up to $53\times$ using GPUs (gpu_dedup, ; storegpu, ) and faster hashing algorithms (tunable_encrypted_dedup, ; persistent_memory_dedup, ), data chunking acceleration has only received limited attention and is currently the performance bottleneck (§3.1).
@@ 2.1. Data Chunking
Data chunking algorithms can generate fixed-size or variable-sized chunks. Dividing the data into fixed-size chunks is fast, but results in poor space savings on most datasets. This is due to the byte-shifting problem (lbfs, ), where adding a single byte causes all subsequent chunks to appear different, despite the data stream largely being unchanged. Thus, while traditional backup systems such as Venti (venti, ) and OceanStore (oceanstore, ) use fixed-size chunks, modern deduplication systems employ content-defined chunking (CDC) algorithms (lbfs, ) to generate variable-sized chunks.
Chunk boundaries in CDC algorithms are derived from the data itself, i.e., they are content-defined. These boundaries are chosen such that most byte shifts cause them to shift by the corresponding amount, leaving subsequent chunks unaffected and preserving the ability to detect duplication. Numerous CDC algorithms have been proposed in previous literature (ae, ; fastcdc, ; gear_hash, ; lbfs, ; maxp, ; ram, ; tttd, ) and can be broadly classified into hash-based and hashless algorithms (dedupbench, ).
@@
These algorithms (fastcdc, ; gear_hash, ; lbfs, ; tttd, ) slide a fixed-size window over the data. When the hash value of the window’s contents matches a target mask, they insert a chunk boundary, creating a new data chunk lying between the current and previous chunk boundaries. Note that these hash-based CDC algorithms are only used during the Data Chunking phase and do not affect the Fingerprinting and Comparison phase.
Algorithms such as Rabin-Karp chunking (lbfs, ) and CRC (sscdc, ) slide a window over the source data and compute the hash of the window’s contents using rolling hash algorithms. Figure 1(a) shows an example of data chunking with such algorithms. In the Rabin-Karp chunking algorithm (lbfs, ), a chunk boundary is declared when the lower k bits of the sliding window’s hash value equals zero. If the current window’s hash value does not meet this condition, the window is slid by a byte. To minimize the overhead of recomputing the hash value, the new value is calculated as a function of the old hash value, the incoming byte, and the outgoing byte (Figure 1(a)), i.e., a rolling hash. This creates a dependency between adjacent bytes, complicating acceleration efforts with SIMD instructions (§3.2). Additionally, despite the development of more lightweight rolling hash algorithms such as CRC (sscdc, ) and Gear-Hash (gear_hash, ), hash-based chunking remains computationally expensive (§6.2).
Some hash-based algorithms like TTTD (tttd, ) and FastCDC (fastcdc, ) use minimum and maximum values to limit the chunk sizes. To improve chunking throughput, these algorithms skip scanning data lying before the minimum chunk size at the beginning of each chunk. Figure 1(b) shows an example of such algorithms with the sub-minimum regions highlighted using a dashed pattern. To offset the impact of skipping the sub-minimum regions and tighten chunk size distributions around the average, FastCDC (fastcdc, ) uses dynamically changing masks, i.e., relaxes the boundary detection condition by reducing k when required. TTTD (tttd, ) uses two different boundary masks to do the same.
@@
AE. Figure 2(a) shows an example chunk generated by the Asymmetric Extremum (AE) (ae, ) algorithm. AE has two modes of operation: AE-Min and AE-Max, depending on whether it uses local minima or maxima; Figure 2(a) shows AE-Max. In each chunk, AE-Max tries to find a target byte that is greater than all the bytes before it. Once the target byte is identified, AE-Max scans a fixed-size window of bytes after the target to identify the maximum-valued byte among them. If the target byte is greater than this maximum-valued byte, it inserts a chunk boundary after the fixed-size window, as shown in Figure 2(a).
@@
MAXP. Figure 2(b) shows an example chunk generated by MAXP (maxp, ). MAXP identifies target bytes in the data stream that are local maxima, i.e., they are greater than a fixed number of bytes before and after them. When such target bytes are found, chunk boundaries are inserted at their locations, as shown in the figure. Note that MAXP’s window sizes are typically 70-80% smaller than AE (ae, ) and RAM (ram, ) to generate the same target average chunk size.
@@
RAM. Figure 2(c) shows an example chunk generated by the Rapid Asymmetric Maximum (RAM) (ram, ) algorithm. RAM begins by scanning a fixed-size window at the beginning of each chunk to find the maximum valued byte (F1 in the figure). It then begins scanning at the first byte outside the window, serially comparing these bytes against this maximum value. A chunk boundary is inserted when the first byte that exceeds or equals the maximum is found, e.g., F3 in Figure 2(c).
@@ 2.2. Deduplication Metrics
Previous literature (understanding_dedup_ratios, ; dedupbench, ; lbfs, ; fastcdc, ) outlines three important metrics for deduplication systems: Space Savings, Chunk Size Distribution, and Chunking Throughput. We describe these in detail in this section.
@@
Space savings (understanding_dedup_ratios, ; dedupbench, ) is one of the primary metrics used to evaluate deduplication systems in production. It represents the overall disk space conserved by using the deduplication system, i.e. size of data stored on disk after deduplication. The space savings achieved are largely dictated by the choice of the data chunking algorithm and its associated parameters. It is defined as:
@@
Chunking throughput is defined as the speed at which the deduplication system divides incoming data into chunks. As CDC algorithms are content-dependent, they need to scan every byte of an incoming data stream before making content-defined boundary decisions. Their speed depends on their computational complexity. Hash-based algorithms involve the use of expensive rolling-hash algorithms to determine chunk boundaries (§2.1), typically resulting in lower throughputs than their hashless counterparts. Chunking throughput is defined as:
@@
Content-defined chunking (CDC) algorithms generate variable-sized chunks of a target average chunk size. They try to ensure that the sizes of generated chunks are as close to the target average as possible. However, due to underlying algorithmic characteristics, each algorithm has a unique chunk size distribution pattern. For instance, FastCDC (fastcdc, ) exhibits two distinct smooth distributions, changing its pattern at the average chunk size. This is because it switches masks past the target average size and relaxes boundary conditions. On the other hand, algorithms such as TTTD (tttd, ) exhibit a smooth distribution between their minimum and maximum specified chunk sizes. Chunk size distributions are typically represented using cumulative distribution function (CDF) plots.
Space savings are inversely proportional to the target average chunk size, i.e., the greater the average chunk size, the lower the space savings achieved in general (primary_dedup, ). This is because the probability of finding duplicate chunks is higher at smaller chunk sizes. The degree of space savings degradation with increasing chunk size depends on algorithmic and dataset characteristics.
All chunks generated by CDC algorithms are subsequently hashed using a collision-resistant algorithm (tunable_encrypted_dedup, ) to generate fingerprints, as described above. The set of unique fingerprints observed thus far is stored in a fingerprint database. New incoming chunks are hashed, and their fingerprints are compared against this database to detect duplicates. Thus, smaller and more numerous chunks result in a larger database; specifically, the fingerprint database size is inversely proportional to the chosen average chunk size. A large number of small chunks can negatively impact system throughput, both due to the increased fingerprint database size and the random data accesses caused by these chunks. Thus, CDC algorithms in production typically target average chunk sizes between 22KB–6464KB.

### v2

@@ Abstract.
Content-defined Chunking (CDC) algorithms dictate the overall space savings that deduplication systems achieve. However, due to their need to scan each file in its entirety, they are slow and often the main performance bottleneck within data deduplication. We present VectorCDC, a method to accelerate hashless CDC algorithms using vector CPU instructions, such as SSE / AVX. We analyzed the state-of-the-art chunking algorithms and discovered that hashless algorithms primarily use two data processing patterns to identify chunk boundaries: Extreme Byte Searches and Range Scans. VectorCDC presents a vector-friendly approach to accelerate these two patterns. Using VectorCDC, we accelerated three state-of-the-art hashless chunking algorithms: RAM, AE, and MAXP. Our evaluation shows that VectorCDC is effective on Intel, AMD, ARM, and IBM CPUs, achieving $8.35\times$$26.2\times$ higher throughput than existing vector-accelerated algorithms, and $15.3\times$$207.2\times$ higher throughput than existing unaccelerated algorithms. VectorCDC achieves this without affecting the deduplication space savings.
@@ 1. Introduction
The amount of data generated and stored on the Internet is growing at an exponential rate (Statista, 2024), and is expected to exceed 180 zettabytes per year in 2025. Storage capacity alone is not well positioned to handle this data influx, with the total installed storage capacity in 2020 only being 6.7 zettabytes (Statista, 2024). Cloud storage providers instead support this data growth using alternatives such as novel storage paradigms (Carlson et al., 2014; Chen et al., 1994), distributed file systems (Shvachko et al., 2010; Weil et al., 2006) and caches (Fitzpatrick, 2004; Bronson et al., 2013), mechanisms such as data deduplication (Meyer and Bolosky, 2012; Xia et al., 2016a), alongside additional investments in data protection (Chen and Zhao, 2012).
@@
Previous studies by Microsoft (El-Shimi et al., 2012) and EMC (Wallace et al., 2012) show that a large amount of redundancy exists in the data stored on the cloud, especially in file system backups (El-Shimi et al., 2012; Meyer and Bolosky, 2012), virtual machine backups (Wallace et al., 2012; Shilane et al., 2012b) and shared documents (Henderson, 2011). The redundancy ratios in these workloads are significant, ranging from $50\%$ to $75\%$. Mechanisms such as data deduplication (Meyer and Bolosky, 2012) and compression (Shilane et al., 2012b) are used to conserve storage space by identifying redundant data, eliminating it, and minimizing its storage impact, thereby improving efficiency and reducing costs.
@@
Data deduplication consists of four phases (Xia et al., 2016a); Data Chunking, Chunk Fingerprinting, Metadata Creation, and Data Storage. Data chunking and chunk fingerprinting are the most computationally intensive (Meyer and Bolosky, 2012; Liu et al., 2023) of these. While chunk fingerprinting has received significant optimization attention, with faster hashing algorithms (Li et al., 2020; Tan et al., 2023) and GPUs (Suttisirikul and Uthayopas, 2012; Al-Kiswany et al., 2008), data chunking optimizations have not kept up (§3.1).
@@
In the data chunking phase, the incoming data is divided into small chunks, typically of size 1641-64 KB. Numerous data chunking algorithms exist in current literature (Zhang et al., 2015; Xia et al., 2016b, 2014; Muthitacharoen et al., 2001; Widodo et al., 2017; Eshghi and Tang, 2005) and can be broadly classified into hash-based and hashless algorithms (Liu et al., 2023). As chunking occurs whenever new data is uploaded, this phase is on the critical path, and directly impacts system performance.
@@
Previous efforts have explored accelerating chunking by using vector instructions. SS-CDC (Ni et al., 2019) uses vector instructions to accelerate hash-based data chunking algorithms, such as Rabin-Karp chunking (Muthitacharoen et al., 2001) and Gear-based chunking (Xia et al., 2014). Unfortunately, this approach only leads to modest improvements in chunking throughput, up to $3.13\times$, as shown in §3.2. Parallelizing hash-based chunking using vector instructions is complicated because these algorithms use the rolling hash of a sliding window of bytes to detect boundaries, inherently creating a computational dependency between adjacent bytes. Consequently, SS-CDC processes different regions of the data in parallel using slow scatter / gather vector instructions, limiting its performance (§3.2).
@@
We posit that hashless chunking algorithms are better candidates for vector acceleration. Although they achieve slightly lower space savings compared to their hash-based counterparts (§6), they are up to $2\times$ faster and use simple mathematical operations (e.g., finding a maximum value) that can be accelerated more efficiently using vector instructions. We analyzed state-of-the-art hashless algorithms to understand their design and identify opportunities to leverage vector instructions. We identified that all state-of-the-art hashless algorithms consist of two processing patterns. The first pattern involves finding local minima or maxima in a data region, which we call Extreme Byte Search, and the second pattern involves scanning a range of bytes to find values that are greater or lesser than a target value, which we call Range Scan. We found that, unlike rolling hash functions, these patterns can be efficiently accelerated using vector instructions.
@@
We implemented VectorCDC using five different vector instruction sets: SSE-128, AVX-256, and AVX-512 on Intel / AMD CPUs; NEON-128 on ARM CPUs; and VSX-128 on IBM Power CPUs. We used VectorCDC to accelerate three state-of-the-art hashless chunking algorithms; RAM (Widodo et al., 2017), AE (Zhang et al., 2015), and MAXP (Bjørner et al., 2010), creating VRAM, VAE, and VMAXP, respectively. We compared the performance of our accelerated algorithms with that of state-of-the-art hash-based algorithms, hashless algorithms, and SS-CDC (Ni et al., 2019) accelerated algorithms using 10 diverse datasets.
Our evaluation (§6) shows that VectorCDC-based algorithms achieve $8.35\times$$26.2\times$ higher chunking throughput than those accelerated with SS-CDC. VRAM, VAE, and VMAXP also achieve $5.51\times$$17.6\times$ higher throughput compared to their unaccelerated hashless counterparts, without affecting deduplication space savings. Furthermore, they achieve $15.3\times$$207.2\times$ higher throughput compared to unaccelerated hash-based algorithms. Finally, VectorCDC-accelerated algorithms retain their performance advantage across all five vector instruction sets.
We have made our code publicly available by integrating it with DedupBench [^1] https://github.com/UWASL/dedup-bench (Liu et al., 2023). Due to the large sizes of our datasets (§6), we were unable to release all of them. Instead, we have publicly released one of our datasets (DEB) on Kaggle [^2] https://www.kaggle.com/datasets/sreeharshau/vm-deb-fast25 (Udayashankar et al., 2025b) and provided detailed descriptions to facilitate the recreation of others for future experiments, similar to previous literature (Xia et al., 2016b; Zhang et al., 2015; Widodo et al., 2017).
The rest of this paper is organized as follows: we discuss relevant background about deduplication and vector instructions in Section 2. Section 3 motivates our work by discussing deduplication performance bottlenecks and the inefficiencies encountered by previous work when accelerating hash-based CDC algorithms. Section 4 outlines VectorCDC’s design while Section 5 discusses implementation challenges across vector instruction sets. Section 6 details our evaluation efforts on diverse datasets. We discuss related work in Section 7 and conclude our paper in Section 8.
@@ 2. Background
Data deduplication consists of four phases (Xia et al., 2016a):
@@
  • Fingerprinting and Comparison: Chunks are hashed using a collision-resistant hashing algorithm such as MurmurHash3 (Li et al., 2020) or SHA-256 (Rachmawati et al., 2018) to generate fingerprints. Fingerprints are compared against those previously seen to identify duplicate chunks.
@@
Data chunking and fingerprinting are typically the most computationally intensive phases in deduplication (Meyer and Bolosky, 2012). While fingerprinting has been accelerated up to $53\times$ using GPUs (Suttisirikul and Uthayopas, 2012; Al-Kiswany et al., 2008) and faster hashing algorithms (Li et al., 2020; Song et al., 2023), data chunking acceleration has only received limited attention and adds significant overhead on the deduplication critical path (§3.1).
@@ 2.1. Data Chunking
Data chunking algorithms can generate fixed-size or variable-sized chunks. Dividing the data into fixed-size chunks is fast, but results in poor space savings on most datasets (§6). This is due to the byte-shifting problem (Muthitacharoen et al., 2001), where adding a single byte causes all subsequent chunks to appear different, despite the data stream largely being unchanged. Thus, while traditional backup systems such as Venti (Quinlan and Dorward, 2002) and OceanStore (Kubiatowicz et al., 2000) use fixed-size chunks, modern deduplication systems employ content-defined chunking (CDC) algorithms (Muthitacharoen et al., 2001) to generate variable-sized chunks.
Chunk boundaries in CDC algorithms are derived from the data itself, i.e., they are content-defined. These boundaries are chosen such that most byte shifts cause them to shift by the corresponding amount, leaving subsequent chunks unaffected and preserving the ability to detect duplication. Numerous CDC algorithms have been proposed in previous literature (Zhang et al., 2015; Xia et al., 2016b, 2014; Muthitacharoen et al., 2001; Bjørner et al., 2010; Widodo et al., 2017; Eshghi and Tang, 2005) and can be broadly classified into hash-based and hashless algorithms (Liu et al., 2023).
@@
These algorithms (Xia et al., 2016b, 2014; Muthitacharoen et al., 2001; Eshghi and Tang, 2005) slide a fixed-size window over the data. When the hash value of the window’s contents matches a target mask, they insert a chunk boundary, creating a new data chunk lying between the current and previous chunk boundaries. Note that these hash-based CDC algorithms are only used during the Data Chunking phase and do not affect the Fingerprinting and Comparison phase.
Algorithms such as Rabin-Karp chunking (Muthitacharoen et al., 2001) and CRC (Ni et al., 2019) slide a window over the source data and compute the hash of the window’s contents using rolling hash algorithms. Figure 1(a) shows an example of data chunking with such algorithms. In the Rabin-Karp chunking algorithm (Muthitacharoen et al., 2001), a chunk boundary is declared when the lower k bits of the sliding window’s hash value equals zero. If the current window’s hash value does not meet this condition, the window is slid by a byte. To minimize the overhead of recomputing the hash value, the new value is calculated as a function of the old hash value, the incoming byte, and the outgoing byte (Figure 1(a)), i.e., a rolling hash. This creates a dependency between adjacent bytes, complicating acceleration efforts with SIMD instructions (§3.2). Additionally, despite the development of more lightweight rolling hash algorithms such as CRC (Ni et al., 2019) and Gear-Hash (Xia et al., 2014), hash-based chunking remains computationally expensive (§6.2).
Some hash-based algorithms like TTTD (Eshghi and Tang, 2005) and FastCDC (Xia et al., 2016b) use minimum and maximum values to limit the chunk sizes. To improve chunking throughput, these algorithms skip scanning data lying before the minimum chunk size at the beginning of each chunk. Figure 1(b) shows an example of such algorithms with the sub-minimum regions highlighted using a dashed pattern. To offset the impact of skipping the sub-minimum regions and tighten chunk size distributions around the average, FastCDC (Xia et al., 2016b) uses dynamically changing masks, i.e., relaxes the boundary detection condition by reducing k when required. TTTD (Eshghi and Tang, 2005) uses two different boundary masks to do the same.
@@
AE. Figure 2(a) shows an example chunk generated by the Asymmetric Extremum (AE) (Zhang et al., 2015) algorithm. AE has two modes of operation: AE-Min and AE-Max, depending on whether it uses local minima or maxima; Figure 2(a) shows AE-Max. In each chunk, AE-Max tries to find a target byte that is greater than all the bytes before it. Once the target byte is identified, AE-Max scans a fixed-size window of bytes after the target to identify the maximum-valued byte among them. If the target byte is greater than this maximum-valued byte, it inserts a chunk boundary after the fixed-size window, as shown in Figure 2(a).
@@
MAXP. Figure 2(b) shows an example chunk generated by MAXP (Bjørner et al., 2010). MAXP identifies target bytes in the data stream that are local maxima, i.e., they are greater than a fixed number of bytes before and after them. When such target bytes are found, chunk boundaries are inserted at their locations, as shown in the figure. Note that MAXP’s window sizes are typically 70-80% smaller than AE (Zhang et al., 2015) and RAM (Widodo et al., 2017) to generate the same target average chunk size. MAXP has also been referred to as Local Maximum Chunking (LMC) in previous literature.
@@
RAM. Figure 2(c) shows an example chunk generated by the Rapid Asymmetric Maximum (RAM) (Widodo et al., 2017) algorithm. RAM begins by scanning a fixed-size window at the beginning of each chunk to find the maximum valued byte (F1 in the figure). It then begins scanning at the first byte outside the window, serially comparing these bytes against this maximum value. A chunk boundary is inserted when the first byte that exceeds or equals the maximum is found, e.g., F3 in Figure 2(c).
@@ 2.2. Deduplication Metrics
Previous literature (Dutch, 2008; Liu et al., 2023; Muthitacharoen et al., 2001; Xia et al., 2016b) outlines three important metrics for deduplication systems: Space Savings, Chunk Size Distribution, and Chunking Throughput. We describe these in detail in this section.
@@
Space savings (Dutch, 2008; Liu et al., 2023) is one of the primary metrics used to evaluate deduplication systems in production. It represents the overall disk space conserved by using the deduplication system, i.e. size of data stored on disk after deduplication. The space savings achieved are largely dictated by the choice of the data chunking algorithm and its associated parameters. It is defined as:
@@
Chunking throughput is defined as the speed at which the deduplication system divides incoming data into chunks. As CDC algorithms are content-dependent, they need to scan every byte of an incoming data stream before making content-defined boundary decisions. Their speed depends on their computational complexity. Hash-based algorithms utilize expensive rolling-hash algorithms to determine chunk boundaries (§2.1), typically resulting in lower throughputs than their hashless counterparts. Chunking throughput is defined as:
@@
Content-defined chunking (CDC) algorithms generate variable-sized chunks of a target average chunk size. They try to ensure that the sizes of generated chunks are as close to the target average as possible. However, due to underlying algorithmic characteristics, each algorithm has a unique chunk size distribution pattern. For instance, FastCDC (Xia et al., 2016b) exhibits two distinct smooth distributions, changing its pattern at the target average chunk size. This is because it switches masks past the target average size and relaxes boundary conditions. On the other hand, algorithms such as TTTD (Eshghi and Tang, 2005) exhibit a smooth distribution between their minimum and maximum specified chunk sizes. Chunk size distributions are typically represented using cumulative distribution function (CDF) plots.
Space savings are inversely proportional to the target average chunk size, i.e., the greater the average chunk size, the lower the space savings achieved in general (El-Shimi et al., 2012). This is because the probability of finding duplicate chunks is higher at smaller chunk sizes. The degree of space savings degradation with increasing chunk size depends on algorithmic and dataset characteristics.
All chunks generated by CDC algorithms are subsequently hashed using a collision-resistant algorithm (Li et al., 2020) to generate fingerprints, as described above. The set of unique fingerprints observed thus far is stored in a fingerprint database. New incoming chunks are hashed, and their fingerprints are compared against this database to detect duplicates. Thus, smaller and more numerous chunks result in a larger database; specifically, the fingerprint database size is inversely proportional to the chosen average chunk size. A large number of small chunks can negatively impact system throughput, both due to the increased fingerprint database size and the random data accesses caused by these chunks. Thus, CDC algorithms in production typically target average chunk sizes between 22KB–6464KB.